A test involving 320 men ages 20 to 24 found that their blood cholesterol levels had a mean of 187 mg/dl and a standard deviation of 44.8 mg/dl.

(a) Determine the z-score, to the nearest hundredth, for one of the men who had a blood cholesterol level of 214 mg/dl.
z214 = ________


(b) The z-score for one man was
−1.51
What was his blood cholesterol level? Round to the nearest hundredth.
= _______ mg/dl

Answer by Theo(13342)

the mean is 187 and the is 44.8.

one of the men had a blood cholesterol level of 214.

the z-score for that individual would be z = (214 - 187) / 44.8 = .6026785714.

round to the nearest hundredth to get a z-score of .60.

the z-score for another man was -1.51.

use the same z-score formula to get:

-1.51 = (x - 187) / 44.8.

solve for x to get:

x = -1.51 * 44.8 + 187 = 119.352.